OpenStudy (anonymous):

Find the second and third term of the sequence -7, , , -22, -27

OpenStudy (anonymous):

@agent0smith

OpenStudy (agent0smith):

Looks like it's just subtracting 5...

OpenStudy (agent0smith):

So subtract 5 to get the second, then do it again to get the third

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

haha well i have another question, so i'll post it

OpenStudy (the_fizicx99):

This looks arithmetic, a_n = a_1 + d(n – 1) a_4 = -7 -5(4 – 1) a_4 = -7 -5(3) a_4 = -7 -15 a_4 = -22 <-- checking (/)_(\) a_5 = -7 -5(4) a_5 = -7 -20 a_5 = -27 So just put a_2 = -7 -5(2-1) a_2 = -7 -5(1) a_2 = -7 -5 a_2 = -12